# SemGauss-SLAM: Dense Semantic Gaussian Splatting SLAM

### v2

@@ 4.3 Ablation Study
We perform ablation study on three scenes of Replica dataset[21] in Tab. 7 to validate the effectiveness of feature-level loss and semantic-informed BA in SemGauss-SLAM. Moreover, ablation of semantic-informed BA component is conducted in Tab. 8.
@@
Table 7: Ablation study of our contributions on three scenes of Replica [21].
@@
Table 8: Ablation study of semantic-informed BA on three scenes of Replica [21]. (w/o semantic) without adding semantic constraint loss $\mathcal{L}_{\text{BA-sem}}$; (w/o RGB and depth) without adding RGB and depth constraint loss $\mathcal{L}_{\text{BA-rgb}}$ and $\mathcal{L}_{\text{BA-depth}}$.
@@ 5 Conclusion
We propose SemGauss-SLAM, a novel dense semantic SLAM system utilizing 3D Gaussian representation that enables dense visual mapping, robust camera tracking, and 3D semantic mapping of the whole scene. We incorporate semantic feature embedding into 3D Gaussian for Gaussian semantic representation, leading to real-time dense semantic mapping. Moreover, we propose feature-level loss for 3D Gaussian scene optimization to achieve accurate semantic representation. In addition, we introduce semantic-informed BA that enables joint optimization of camera poses and 3D Gaussian representation by establishing multi-view semantic constraints, resulting in low-drift tracking and precise mapping.

### v3

@@ 4.3 Ablation Study
Figure 5: Semantic rendering results and ground truth labels of feature-level loss ablation on two scenes of Replica straub2019replica .
We perform ablation study on three scenes of Replica dataset straub2019replica in Tab. 5 to validate the effectiveness of feature-level loss and semantic-informed BA in SemGauss-SLAM. Moreover, ablation of semantic-informed BA component is conducted in Tab. 6.
@@
Table 5: Ablation study of our contributions on three scenes of Replica straub2019replica .
@@
Table 6: Ablation study of semantic-informed BA on three scenes of Replica straub2019replica . (w/o semantic) without adding semantic constraint loss $\mathcal{L}_{\text{BA-sem}}$; (w/o RGB and depth) without adding RGB and depth constraint loss $\mathcal{L}_{\text{BA-rgb}}$ and $\mathcal{L}_{\text{BA-depth}}$.
@@ 5 Conclusion
We propose SemGauss-SLAM, a novel dense semantic SLAM system utilizing 3D Gaussian representation that enables dense visual mapping, robust camera tracking, and 3D semantic mapping of the whole scene. We incorporate semantic feature embedding into 3D Gaussian for Gaussian semantic representation for dense semantic mapping. Moreover, we propose feature-level loss for 3D Gaussian scene optimization to achieve accurate semantic representation. In addition, we introduce semantic-informed BA that enables joint optimization of camera poses and 3D Gaussian representation by establishing multi-view semantic constraints, resulting in low-drift tracking and precise mapping.